OpenStudy (anonymous):

Let logbA=3 logbC=2 logbD=5 then what is the value of logb d^2/ c^3a

OpenStudy (anonymous):

\[\log_{b} x = y\] means \[b^y = x\] thus \[\log_{b} A = 3\] \[b^3 = A\] \[\log_{b} C = 2\] \[b^2 = C\] \[\log_{b} D = 5\] \[b^5 = D\] \[\frac{ \log_{b}D^2 }{ c^3*a } = \log_{b} \frac{ (b^5)^2 }{ (b^2)^3 *b^3 }\] so what will be the answer?

OpenStudy (anonymous):

A. 19 B. 4.44 C. 1 D. 0.549

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

\[lob_b \frac{ b^{10} }{ b^9 }\] \[\log_b b = 1\]